Suppose J is between H and K. Use the segment addition postulate to solve for x. Then find the length of each segment.
The problem is: HJ=3(x+2)
JK=3x-4
KH=44

Found 2 solutions by palanisamy, checkley77:
Answer by palanisamy(496)

HJ+JK = KH
3(x+2)+(3x-4) = 44
3x+6+3x-4 = 44
6x+2 = 44
6x=44-2
6x=42
x = 42/6
x = 7

Answer by checkley77(12844)

3(x+2)+3x-4=44
3x+6+3x-4=44
6x=44-6+4
6x=42
x=42/6
x=7 ans.
Proof:
3(7+2)+3*7-4=44
3*9+21-4=44
27+21-4=44
44=44